OpenStudy (anonymous):

find the missing term 2,3,6,15,X,157.5,630

OpenStudy (tkhunny):

Fill in the value with a term you can justify. There are infinitely many. Just pick one and justify it. I REALLY HATE these poorly worded problems. I'll probably just give you an answer most of the time. 2/3 = 1.5 6/3 = 2.0 15/6 = 2.5 X/15 = ?? 157.5/X = ??? 630/157.5 = 4 I'm seeing something. That right-hand side might go like this: 1.5, 2.0, 2.5, 3.0, 3.5, 4.0 Give it a try.